OpenStudy (adalrymple):

HELP PLEASE! solve the system of linear equations: 3x+4y=18 x+y=11

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

the second equation is the same as \[y=-x+11\] put that in the first and solve \[3x+4(-x+11)=18\] for \(x\)

OpenStudy (adalrymple):

Im confused.. it doesn't work.

OpenStudy (gnorris):

always helps

OpenStudy (karimsqualli96):

@adalrymple you have a system of two equations with two unknows, to find the solution you must find pair (x,y) that solve the system( when you put (x,y) in each equation, its give the right result) here there is unique solution to find that solution : in the second equation move x to the right side , this give : y = -x + 11; know take that y and replace it in the first equation : 3x + 4y = 18 3x + 4(-x + 11) = 18 3x -4x +44 = 18 -x = 18 - 44 x = 26 we have x know, to find y replace in : y = -x + 11 y = -26 + 11 y = -15

OpenStudy (adalrymple):

@karimsqualli96 thank you :)

OpenStudy (karimsqualli96):

@adalrymple you're welcome :)